OpenStudy (prince1342):

MEDAL AND FAN 4. The theme park company is building a scale model of the Killer whale stadium main show tank for an investor's presentation. How many times smaller will the Volume of the model be if each dimension were made 6 times smaller?

OpenStudy (prince1342):

@KJSaif

OpenStudy (mudkillz):

the volume would be 6 time smaller.

OpenStudy (mudkillz):

@Prince

OpenStudy (prince1342):

so i divide it by 6?

OpenStudy (mudkillz):

Yes, that is what you would do

OpenStudy (mudkillz):

Since the numbers used to make the calculation have become 6 times smaller, so would the answer

OpenStudy (prince1342):

Would I divide 6 by 0.03590065 or 0.03590065 by 6?

OpenStudy (jdoe0001):

 \(\bf volume=length\cdot width\cdot height \\ \quad \\ \frac{1}{6}\textit{ of each of those units, means} \\ \quad \\ \cfrac{length}{6}\cdot \cfrac{width}{6}\cdot \cfrac{height}{6}\implies \cfrac{length\cdot width\cdot height}{6\cdot 6\cdot 6} \\ \quad \\ \cfrac{length\cdot width\cdot height}{6^3}\implies \cfrac{volume}{6^3}\)